cos (2001)pi + cot(2001)pi/2 + sec (2001) pi/3 + tan(2001)pi/4 + cosec(2001) pi/6 + 2 is equal to
 

Grade:11

5 Answers

avanthikaa sri
6 years ago
cos(2001)pi is in the form of cosn.pi=-1
cot(2001)pi/2=0
sec(2001)pi/3=sec667.pi/3=-1
tan(2001)pi/4=tan(500pi+pi/4)=1
cosec(2001)pi/6=-1
-1+0-1+1-1+2
-2+2
0 is the answer
mhb
2 years ago
os(2001)pi is in the form of cosn.pi=-1
cot(2001)pi/2=0
sec(2001)pi/3=sec667.pi/3=-1
tan(2001)pi/4=tan(500pi+pi/4)=1
cosec(2001)pi/6=-1
-1+0-1+1-1+2
-2+2
0 is the answer
mhb
2 years ago
os(2001)pi is in the form of cosn.pi=-1
cot(2001)pi/2=0
sec(2001)pi/3=sec667.pi/3=-1
tan(2001)pi/4=tan(500pi+pi/4)=1
cosec(2001)pi/6=-1
-1+0-1+1-1+2
-2+2
 the answer
mhb
2 years ago
os(2001)pi is in the form of cosn.pi=-1
cot(2001)pi/2=0
sec(2001)pi/3=sec667.pi/3=-1
tan(2001)pi/4=tan(500pi+pi/4)=1
cosec(2001)pi/6=-1
-1+0-1+1-1+2
-2+2
 
 
mhb
2 years ago
os(2001)pi is in the form of cosn.pi=-1
cot(2001)pi/2=0
sec(2001)pi/3=sec667.pi/3=-1
tan(2001)pi/4=tan(500pi+pi/4)=1
cosec(2001)pi/6=-1
-1+0-1+1-1+2